OpenStudy (anonymous):

June Davison borrowed $9,000 for 31/2 years at 12% simple interest. Find the interest

OpenStudy (anonymous):

when we are just told interest, we assume it is yearly. simple interest is different from compound interest. simple interest is always calculated from the initial borrowed amount every year whereas compound is calculated from the updated amount that gets larger with every interest calculation [at the end of every year]. I = PV*i*N I = Interst PV = Present value i = interest rate [yearly] N = amount of years I = 9000(0.12)(3.5) = $3780

OpenStudy (anonymous):

Thank you for the really good explanation. :D

OpenStudy (anonymous):

glad i could help ^_^